A hamburger wrapper is thrown from a third floor dorm window, and its motion can be modeled by the equation h(t) = -3t^2 + 6t + 30, where h is the height in feet and t is the time in seconds.


a. Find the vertex of the parabola, then describe what it tells us physically about the situation.

b. Find the x-intercept, then describe what it tells us physically about the situation.





1
Expert's answer
2021-12-02T12:17:53-0500

"h(t)=-3t^2+6t+30"


a. "a=-3,\ \ b=6,\ \ t=-\frac{b}{2a}=-\frac{6}{-6}=1"

"h(1)=-3+6+30=33"

Vertex of parabola: "t=1,\ \ h=33" .

It means that the hamburger wrapper will reach the maximum height "h=33" feet at the time "t=1" second.


b. "h(t)=0,\ \" "a=-3,\ \ b=6,\ \ c=30"

"t_{1,2}=\frac{-b\pm\sqrt{b^2-4ac}}{2a}=\frac{-6\pm\sqrt{36+360}}{-6}=\frac{-6\pm 6\sqrt{11}}{-6}=1\pm\sqrt{11}"


"t_1=1-\sqrt{11}<0" ("t" is the time in seconds, so "t\geq 0" )

"t_2=1+\sqrt{11}\approx 4.3" seconds is the time when the hamburger wrapper reaches the ground (flight time)


Answers: a. "(1;\ 33)" , b. "t=1+\sqrt{11}" (and "1-\sqrt{11}" )